OpenStudy (errinss):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 256 A. 128 B. 128, –128 C. 16, –16 D. 16

OpenStudy (cheez):

I am guessing this is x^2 and not 2x. sqrtx^2 = sqrt256 x = 16.

OpenStudy (cheez):

BUT it can be both negative and positive, so C.

OpenStudy (errinss):

ya i meant to make it the 2 an exponent

OpenStudy (errinss):

Solve the equation. If necessary, round your answer to the nearest tenth. 4x2 = 81 A. 2.25, –2.25 B. 4.5, –4.5 C. 9, –9 D. 20.25, –20.25

OpenStudy (errinss):

\[4x^{2}=81\]

OpenStudy (cheez):

First, for your second one, try dividing both sides by 4. what do you get? (imma let you try to do this one on ur own.)

OpenStudy (errinss):

i got 20.25

OpenStudy (cheez):

good. Now, square root both sides again. You should get--- x = ? Find ?

OpenStudy (errinss):

4.5?

OpenStudy (cheez):

yes! Great job! but it can be both negative or positive, so C. again.

OpenStudy (cheez):

I meant B. :P

OpenStudy (cheez):

need any more help?

OpenStudy (errinss):

ya

OpenStudy (cheez):

Ok, shoot

OpenStudy (errinss):

What is the value of x? Round your answer to the nearest tenth. x^2 = 112

OpenStudy (cheez):

Easy! do the same as the first one, sqrt both sides. What do you get?

OpenStudy (errinss):

10.583005244258362362006463014557 this is what i got...

OpenStudy (cheez):

Ok so now round it up. 8 is greater than 5, so it becomes 10.6

OpenStudy (errinss):

but wouldn't you round in the tenths place instead of the hundredths place

OpenStudy (cheez):

and no it's not is it greater than the number before it. if you have at least .5 behind a number, you round it up. For example: 6.26 because 6 is greater than 5, you round it up to 6.3

OpenStudy (cheez):

10.6 is the tenths place. x = a number. .x = tenths .xx = hundredths .xxx = thousandths and so on.

OpenStudy (errinss):

oh that makes sense lol

OpenStudy (cheez):

Hope I helped!

OpenStudy (cheez):

If you ever need help just tag me. y'know with the @ thing.

OpenStudy (errinss):

ya can you help me with 1 more?

OpenStudy (errinss):

\[4y^{2}+8=264\]

OpenStudy (errinss):

what is y?

pooja195 (pooja195):

can you make a new post?

OpenStudy (errinss):

ya

OpenStudy (cheez):

Sorry I was AFK for a sec. @errinss make the X all by itself. by that I mean subtract 8 from both sides. Thus, you get 4y^2 = 256 Divide both sides by 4, then square root both sides of the new equation. you should get y = 8.